# WildRefer: 3D Object Localization in Large-scale Dynamic Scenes with Multi-modal Visual Data and Natural Language

### v2

@@ 1 Introduction
To facilitate the research of 3DVGW, we propose two new datasets, STRefer and LifeRefer, which are captured in large-scale human-centric scenarios by an 128-beam mechanical LiDAR and an industrial cameras. STRefer is extended from the open-source dataset STCrowd [12], which targets for 3D perception task in crowd scenarios. We sample 662 LiDAR scans and synchronized images and add 5,458 free-form language descriptions. LifeRefer is extended from the huge dataset HuCenLife [47], collected in various daily-life scenes with rich human activities, human-human interactions, and human-object interactions. We annotate 25,380 natural language descriptions for 11,864 subjects from 3,172 scenes. Human-centric free environment is much more complex than static indoor scenes or traffic scenes, which is challenging but significant for the development of autonomous driving and service robots. Our datasets fill the blank of data in this field and have potential to boost related research. In particular, our method is evaluated on both datasets by extensive experiments and achieve state-of-the-art performance.
@@ 4.2 Feature extraction
We use Pointnet++ [39] to extract the geometry features $\{F_{P_{t}},F_{P_{t-1}},F_{P_{t-2}},...,F_{P_{t-K+1}}\}$ from temporal point clouds and pretrained ResNet34 [18] to extract appearance features $\{F_{I_{t}},F_{I_{t-1}},F_{I_{t-2}},...,F_{I_{t-K+1}}\}$ from temporal images, where subscripts $t$ and $t-1$ denote the data of current and previous time, and $K$ denotes the length of input sequence. For utterances, we leverage pretrained RoBERTa [36] to extract text feature $F_{L}$.
Figure 3: Structure of dynamic visual encoder. $F$ indicates the feature extracted from either LiDAR points cloud or images, $t$ denotes the time, and PosEmb means position embedding.
@@ 4.3 Dynamic Visual Encoder (DVE)
Apart from location or texture related descriptions, dynamic motion-related descriptions are also important signals to locate specific human target in large-scale scenes, such as walking, standing, sitting, riding, waving, etc. Moreover, the occlusion problem in dense crowd can also be alleviated by using temporal information. Such dynamic features can be extracted from consecutive visual data, where the changes of depth values in points clouds and pixel values in images represent motion patterns of human actions. Because the scene content is always changing, it is not easy to match the object features from different times in large-scale scenes. Therefore, we apply a spatial-temporal attention mechanism to track critical features for objects automatically. As shown in Figure. 3, we first enhance the spatial feature of current time by self-attention. And then, by gradually matching the feature from previous time series through cross-attention, we obtain the dynamic feature $F$ of object. The process for utilizing two consecutive frames is
@@ 5.3 Ablation Studies
Figure 5: Visualization results of ablation results on STRefer.
@@ 5.5 Temporal Information Exploration
6 Future Work
Our new datasets all focus on human targets in large-scale scenarios, which is significant for the development of service robots and intelligent security. However, it is promising to extend them with more comprehensive descriptions for more kinds of objects in scenes. By then, human-object interactions and human-environment interactions should be considered for more accurate 3D localization.
@@ 7 Conclusion
In this paper, we first introduce the task of 3DVGW, which focuses on the large-scale dynamic scenes and uses online captured multi-modal visual data. Then, we propose effective vision-language interaction and sensor-fusion strategy for 3D vision grounding by fully utilizing the geometry feature, appearance feature, and motion feature. In particular, we propose two novel datasets for 3DVGW on human-centric scenarios and hope our work could inspire more related research for developing AI robots.

### v3

@@ 1 Introduction
To facilitate the research of 3DVGW, we propose two new datasets, STRefer and LifeRefer, which are captured in large-scale human-centric scenarios by a 128-beam mechanical LiDAR and an industrial camera. STRefer is extended from the open-source dataset STCrowd [12], which targets for 3D perception task in crowd scenarios. We sample 662 LiDAR scans and synchronized images and add 5,458 free-form language descriptions. LifeRefer is extended from the huge dataset HuCenLife [47], collected in various daily-life scenes with rich human activities, human-human interactions, and human-object interactions. We annotate 25,380 natural language descriptions for 11,864 subjects from 3,172 scenes. Human-centric free environment is much more complex than static indoor scenes or traffic scenes, which is challenging but significant for the development of autonomous driving and service robots. Our datasets fill the blank of data in this field and have potential to boost related research. In particular, our method is evaluated on both datasets and achieves state-of-the-art performance.
@@ 4.2 Feature extraction
We use Pointnet++ [39] to extract the geometry features $\{F_{P_{t}},F_{P_{t-1}},...,F_{P_{t-K+1}}\}$ from temporal point clouds and pretrained ResNet34 [18] to extract appearance features $\{F_{I_{t}},F_{I_{t-1}},...,F_{I_{t-K+1}}\}$ from temporal images, where $t$ denotes current time, and $K$ denotes the length of input sequence. For utterances, we leverage pretrained RoBERTa [36] to extract text feature $F_{L}$.
Figure 3: Structure of dynamic visual encoder. $F$ indicates the feature extracted from either point clouds or images, $t$ is the time, and PosEmb means position embedding.
@@ 4.3 Dynamic Visual Encoder (DVE)
Apart from location or texture related descriptions, dynamic motion-related descriptions are also important signals to locate specific human target in large-scale scenes, such as walking, standing, sitting, riding, waving, etc. Moreover, the occlusion problem in dense crowd can also be alleviated by using temporal information. Such dynamic features can be extracted from consecutive visual data, where the changes of depth values in points clouds and pixel values in images represent motion patterns of human actions. Because the scene content is always changing, it is not easy to match the object features from different times in large-scale scenes. Therefore, we apply a spatial-temporal attention mechanism to track critical features of objects automatically. As Figure. 3 shows, we first enhance the spatial feature of current time by self-attention. Then, by gradually matching the feature in previous time series through cross-attention, we obtain the dynamic feature $F$ of object. The process of using two adjacent frames is
@@ 5.2 Comparison
Figure 5: Visualization results of ablation results on STRefer.
@@ 6 Conclusion
In this paper, we first introduce the task of 3DVGW, which focuses on the large-scale dynamic scenes and uses online captured multi-modal visual data. Then, we propose effective vision-language interaction and sensor-fusion strategy for 3D vision grounding by fully utilizing the geometry feature, appearance feature, and motion feature. In particular, we propose two novel datasets for 3DVGW on human-centric scenarios and hope our work could inspire more related research for developing AI robots. In the future, it is promising to extend them with more comprehensive descriptions for more kinds of objects in scenes.